# A Bird Song Detector for improving bird identification through Deep Learning: a case study from Doñana

### v1

@@ Abstract
Passive Acoustic Monitoring, which uses devices like automatic audio recorders, has become a fundamental tool in conserving and managing natural ecosystems. However, the large volume of unsupervised audio data it generates poses a major challenge for extracting meaningful information. Deep Learning techniques, particularly automated species identification models based on computer vision, offer a promising solution. BirdNET, a widely used model for bird identification, has shown success in many study systems but is limited in some regions due to biases in its training data, which focus on specific locations and target sounds rather than entire soundscapes. A key challenge in bird species detection is that many recordings either lack target species or contain overlapping vocalizations, complicating automatic identification. To overcome these problems, we developed a multi-stage pipeline for automatic bird vocalization identification in Doñana National Park (SW Spain), a region facing significant conservation threats. Our working pipeline included, first, the development of a Bird Song Detector to isolate bird vocalizations, and second, custom classifiers trained with BirdNET’s embeddings for species classification at the local scale. We manually annotated 461 minutes of audio data from three main habitats across nine different locations within Doñana, resulting in 3749 annotations covering 34 classes. Spectrograms were employed as graphical representations of bird audio data, facilitating the application of image processing methods. The best-performing detection model incorporated synthetic background audios through data augmentation and an environmental sound library (ESC-50). Applying the Bird Song Detector before classification improved species identification, as all classification models performed better when analyzing only the segments where birds were detected. Specifically, the combination of the Bird Song Detector and fine-tuned BirdNET achieved increases in weighted precision (from 0.18 to 0.37), recall (from 0.21 to 0.30), and F1 score (from 0.17 to 0.28), compared to the baseline without the Bird Song Detector. Our approach demonstrated the effectiveness of integrating a Bird Song Detector with fine-tuned classification models for bird identification at local soundscapes. These findings highlight the need to adapt general-purpose tools for specific ecological challenges, as demonstrated in Doñana. Automatically detecting bird species serves for tracking the health status of this threatened ecosystem, given the sensitivity of birds to environmental changes, and helps in the design of conservation measures for reducing biodiversity loss.
@@
Computer Vision , Convolutional Neural Networks , Deep Learning , Ecoacoustics , Passive Acoustic Monitoring
@@ 1 Introduction
In the case of bird vocalization recognition, Deep Learning techniques have represented a revolution for monitoring bird populations (Xie et al., 2023). One of the most popular models of bird vocalization recognition is BirdNET (Kahl et al., 2021), which has proved to be successful in many cases (Pérez-Granados, 2023; Wood and Kahl, 2024; Schuster et al., 2024). These models were tested in environments where the base model was especially well trained, mainly in Northern America and central-Northern Europe. This means that the model was initially trained on a dataset that closely resembles the conditions of the test environment, thereby increasing its predictive accuracy. Nevertheless, models trained on global datasets performance may degrade when applied to unfamiliar environmental conditions, given the variability of local bird vocalizations and soundscapes (Beery et al., 2019; Lauha et al., 2022). While BirdNET performs well in regions it was trained on, its accuracy declines in unfamiliar soundscapes due to local variations in bird vocalizations, background noise, and overlapping bird vocalizations (Pérez-Granados, 2023). In these cases, False Positives (FPs) often arise from other vocalizing animals not being birds, anthropogenic sounds, or weather conditions (Stowell et al., 2019; Kahl et al., 2021; Clark et al., 2023).
@@
Foundational models like BirdNET also struggle to recognize species they were not trained on. In theory, it is possible to retrain an existing model to add missing species, known as fine-tuning (Lalor et al., 2017). Alternatives such as Google’s Perch model (Hamer et al., 2023) and custom models based on BirdNET with fine-tuning for local conditions have been developed (Brunk et al., 2023; Sossover et al., 2024; Ghani et al., 2024). These fine-tuned models aim to improve accuracy by accounting for the unique characteristics of local bird populations (Lauha et al., 2022). However, this task is quite challenging as it requires Machine Learning expertise similar to having to train models from scratch. While recent studies have demonstrated progress in applying Machine Learning to accelerate the annotation process in ecoacoustics (Martin et al., 2022; Sethi et al., 2024), these approaches are still under development and not yet widely implemented.
@@
To address these limitations, we have been inspired by methodologies used in camera trap projects (Beery et al., 2019; Rigoudy et al., 2023), in which a detector is applied prior to species classification on captured images. Similarly, we propose a multi-stage pipeline approach that uses first a generalizable bird vocalization detector combined later with a classifier, improving accuracy and reducing false positives. In particular, we have developed a Bird Song Detector based on the YOLOv8 model (You Only Look Once v8; Jocher et al. (2023)) trained with data from our case study in Doñana National Park. Once the audio segments containing bird songs have been detected, we applied several classifier models, and in all of them, we have proven an improvement of the classification when the detector was present (Figure 1).
@@ 2.1.1 Field study site and PAM design
Soundscapes were recorded at Doñana National Park (SW Spain). This area corresponds to the marshes of the Guadalquivir delta and is one of the most important wetlands in southern Europe, where millions of migrating birds stopover and winter every year (Rendón et al., 2008; Green et al., 2016). The area has serious conservation threats due to water over-exploitation for agriculture and tourism, climate and land-use change, and pollution due to a mine spillover in 1998, among others (Green et al., 2024), which poses serious concerns both at the local and European administration levels. These threats are particularly affecting bird populations, which are in decline over the last years (Camacho et al., 2022; Campo-Celada et al., 2022). Doñana includes four major habitats, which are differentiated by their flooding regime and vegetation: coastal dunes, scrublands, marshlands, and the ecotone or transition among them. The deployment design of the BIRDeep project included nine AudioMoth recorders (Hill et al., 2019) that were distributed among these three habitats: two in the marshland, three in the ecotone and four in the scrubland, differentiating high and low scrubland (see Figure 2). AudioMoths are low-cost automatic audio recording devices with open-source hardware (Hill et al., 2018). They continuously recorded 1 minute of audio every 10 minutes. Configuration parameters of deployed AudioMoth included a sampling rate of 32 kHz, a medium gain, and a filter band focused on bird frequencies (0.6-16.0 kHz).
@@ 2.1.2 Acoustic data annotation by experts
A total of 461 minutes of audio data, recorded between March and October 2023, were annotated by two of the co-authors of this work with ornithological expertise (ESG and GB). They used Audacity software to listen to the recordings and annotate the spectrograms (Audacity, 2017). Annotations were exported in Audacity format and then converted to CSVs. Annotation efforts prioritized periods of high bird activity, mainly at morning chorus, to maximize the number of bird vocalizations (Robbins, 1981). Each annotation consisted of bounding boxes with the minimum and maximum frequency and start time and end time of each bird vocalization in each spectrogram.
@@
The number of annotations was highly unbalanced across recorders, whereas the number of annotated files was more similar, except for Juncabalejo site (Figure 3). Some recorders had a high number of annotations because their recordings were rich in bird vocalizations, likely reflecting variations in bird occurrence and activity, environmental conditions, and recording quality across sampling sites.
@@
Figure 3: Frequency distribution of annotation and audio file across sampling sites (see Figure 2 for reference of site names), illustrating the varying levels of annotation effort across recorders.
@@
Figure 4: Frequency distribution of the 34 annotated classes in the dataset. All classes have been transformed to the Bird class to train the Bird Song Detector. Bars in orange indicate classes that are removed to train the classifiers (29 classes).
@@ 2.1.4 Data Preparation
Spectrograms were generated using the librosa Python library (McFee et al., 2015), a popular toolkit for audio analysis. The Short-Time Fourier Transform (STFT) was applied to the audio signal to compute the spectrogram, and the amplitude of the resulting frequency bins was converted to a decibel (dB) scale. The frequency axis was displayed on a logarithmic scale. The frequency range of the spectrograms was set to a minimum of 1 Hz and a maximum of 16,000 Hz, which encompasses the typical range of bird vocalizations while excluding inaudible frequencies or irrelevant noise. The dimensions of the output spectrogram images were $930\times 462$ pixels. Although our annotations originally included frequency windows for each vocalization, we simplified this by considering the full frequency spectrum of each vocalization to simplify processing, given the limited size of the dataset (see Figure 6 as an example).
@@
To enhance the dataset and improve the robustness of downstream models, additional preprocessing steps were applied. Specifically, synthetic audio samples were generated by adding random noise (using NumPy Python library (Harris et al., 2020), to simulate background interference and slightly adjusting the intensity of the signals. These modifications aimed to increase the number of training examples that contained only background noise, ensuring a more diverse dataset for both detection and classification tasks.
The annotations were further processed to adopt the format required by the YOLOv8-based object detection model (Jocher et al., 2023). Coordinates must be relative to the image dimensions and normalized. Each annotation was then converted into bounding boxes, where temporal coordinates were calculated based on the start and end times of the annotation, the duration of the audio segment (60 seconds) and the width of the spectrogram (930 pixels). For the frequency axis, the bounding box was centered along the y-axis of the spectrogram and spanned the entire height, represented as y_center = 0.5 and y_height = 1 in normalized terms. This process resulted in bounding box values represented as x_center, y_center, x_width, y_height, normalized between 0 and 1. In line with YOLO’s official documentation (Ultralytics, 2024), annotations were stored in individual *.txt files for each spectrogram in a mirror directory corresponding to the audio files. The format of these files includes one row per annotation: class, x_center, y_center, x_width, y_height. Each class is zero-indexed (starting from 0). Spectrograms without annotations have an empty *.txt file.
@@ 2.2 Bird Song Detector Model Development
2.2.1 Bird Song Detector Predictions
The Bird Song Detector’s predictions are represented as bounding boxes in the YOLOv8 output format. Each bounding box is defined by the image-normalized coordinates of its center ($x_{\text{center}},y_{\text{center}}$), along with its normalized width ($w$) and height ($h$). To convert these image-based coordinates into temporal coordinates, we use the $x_{\text{center}}$ value and the image width $W$ to determine the start and end times of the detected bird song segment. This transformation follows these equations:
$x_{\textit{center\_d}}=x_{\textit{center}}\times\textit{W}$
[…]
where $W$ is the width of the spectrogram image used as input for the Bird Song Detector, set to 930 pixels in our case. The variables $x_{\text{center}}$ and $w$ are normalized values representing the bounding box center and width, respectively. To map these predictions to real-world time, they are first denormalized into $x_{\textit{center\_d}}$ and $w_{\textit{d}}$ by scaling with $W$. The resulting values, start_sec and end_sec, correspond to the start and end times (in seconds) of the detected bird song segment relative to that segment.
@@ 2.2.2.2   BirdNET as a bird vocalization detector
    0.1 Confidence Score (Default Value): This is the default BirdNET threshold. Prior studies have shown that using this threshold in real-world PAM scenarios leads to an increase in FPs, as BirdNET assigns low-confidence scores to background noise or ambiguous detections (Pérez-Granados, 2023). However, lowering the confidence threshold improves recall by reducing missed detections.
@@
    0.6 Confidence Score (Adjusted Threshold): Higher confidence thresholds have been recommended for improving the reliability of BirdNET predictions in complex acoustic environments. Funosas et al. (2024) found that BirdNET predictions become significantly more reliable when using thresholds above 0.7. We adopted a slightly lower threshold of 0.6 to balance precision and recall, reducing FPs while still capturing a sufficient number of detections.
@@ 2.3.2 Other classifiers
During evaluation, the models were tested using full 1-minute audio files segmented into 3-second windows and also using only the detections from the Bird Song Detector, padding shorter detections to fit the 3-second input size.
@@ 3.2 Selection of confidence score threshold for the Bird Song Detector
On the other hand, higher thresholds, like 80% or 95%, lead to excessive losses of TPs (74.35% and 99.03%, respectively). These thresholds result in an impractical trade-off, where the reduction in FPs comes at the cost of almost complete loss of the TPs (which need to have the high probability threshold to be retained), significantly degrading the detector’s performance.
@@ 3.3 Bird Song Detector Selection
In the case of the Full Frequencies model, although higher precision is achieved, the limited number of predictions raises concerns about its robustness and generalizability. In contrast, the Best Model, with a greater number of predictions and balanced performance metrics, stands out as the more reliable choice for our ecological monitoring objectives given its consistent performance across metrics.
@@ 4.2 Bird Song Detector vs BirdNET as a Bird Vocalization Detector
Table 2: Comparison of confusion matrices for different confidence score thresholds and species lists in BirdNET (non fine-tuned and fine-tuned) and Bird Song Detector model. The table shows the number of TPs, FPs and FNs for each configuration. The last two rows highlight the relative performance changes when comparing the Bird Song Detector with Fine-tuned BirdNET with different confidence thresholds (in blue, 0.6 and in orange, 0.1), focusing on the detection accuracy for bird sounds versus background noise. Bold values indicate improvements made by Bird Song Detector.
@@ 4.3 Bird Species Classification
Table 3: Comparison of classifier performance with and without the Bird Song Detector. Shown metrics are Accuracy (Acc.), macro and weighted average (Avg) precision (Prec.), recall (Rec.), F1-score (F1), and an index calculated based on the number of predictions relative to the total number of annotations (Idx Pred/Annn). Shaded rows indicate improvement when using Bird Song Detector.

### v2

@@ Abstract
Passive Acoustic Monitoring (PAM), which uses devices like automatic audio recorders, has become a fundamental tool in conserving and managing natural ecosystems. However, the large volume of unsupervised audio data that PAM generates poses a major challenge for extracting meaningful information. Deep Learning techniques, particularly automated species identification models based on computer vision, offer a promising solution. BirdNET, a widely used model for bird identification, has shown success in many study systems but is limited at local scale due to biases in its training data, which focus on specific locations and target sounds rather than entire soundscapes. A key challenge in bird species detection is that many recordings either lack target species or contain overlapping vocalizations, complicating automatic identification. To overcome these problems, we developed a three-stage pipeline for automatic bird vocalization identification in Doñana National Park (SW Spain), a wetland facing significant conservation threats. We deployed AudioMoth recorders in three main habitats across nine different locations within Doñana, and the manual annotation of 461 minutes of audio data, resulting in 3749 annotations covering 34 classes. Our working pipeline included, first, the development of a Bird Song Detector to isolate bird vocalizations, using spectrograms as graphical representations of bird audio data and applying image processing methods. Second, we classified bird species training custom classifiers at the local scale with BirdNET’s embeddings. The best-performing detection model incorporated synthetic background audios through data augmentation and an environmental sound library (ESC-50). Applying the Bird Song Detector before classification improved species identification, as all classification models performed better when analyzing only the segments where birds were detected. Specifically, the combination of the Bird Song Detector and fine-tuned BirdNET increased weighted precision (from 0.18 to 0.37), recall (from 0.21 to 0.30), and F1 score (from 0.17 to 0.28), compared to the baseline without the Bird Song Detector. Our approach demonstrated the effectiveness of integrating a Bird Song Detector with fine-tuned classification models for bird identification at local soundscapes. These findings highlight the need to adapt general-purpose tools for specific ecological challenges, as demonstrated in Doñana. Automatically detecting bird species serves for tracking the health status of this threatened ecosystem, given the sensitivity of birds to environmental changes, and helps in the design of conservation measures for reducing biodiversity loss.
@@
AudioMoth , Computer Vision , Convolutional Neural Networks , Ecoacoustics , Ecosystem health , Passive Acoustic Monitoring
journal: Ecological Informatics
@@ 1 Introduction
In the case of bird vocalization recognition, Deep Learning techniques have represented a revolution for monitoring bird populations (Xie et al., 2023; Stowell, 2022). One of the most popular models of bird vocalization recognition is BirdNET (Kahl et al., 2021), which has proved to be successful in many cases (Pérez-Granados, 2023; Wood and Kahl, 2024; Schuster et al., 2024). These models were tested in environments where the base model was especially well trained, mainly in Northern America and central-Northern Europe. This means that the model was initially trained on a dataset that closely resembles the conditions of the test environment, thereby increasing its predictive accuracy. While BirdNET performs well in regions it was trained on, its accuracy declines in unfamiliar soundscapes due to local variations in bird vocalizations, background noise, and overlapping bird vocalizations (Beery et al., 2019; Lauha et al., 2022; Pérez-Granados, 2023). In these cases, False Positives (FPs) often arise from other vocalizing animals not being birds, anthropogenic sounds, or weather conditions (Stowell et al., 2019; Kahl et al., 2021; Clark et al., 2023).
@@
Foundational models like BirdNET also struggle to recognize species they were not trained on. In theory, it is possible to retrain an existing model to add missing species, known as fine-tuning (Lalor et al., 2017). Alternatives such as Google’s Perch model (Hamer et al., 2023) and custom models based on BirdNET with fine-tuning for local conditions have been developed (Brunk et al., 2023; Sossover et al., 2024; Ghani et al., 2024; Pérez-Granados et al., 2025). These fine-tuned models aim to improve accuracy by accounting for the unique characteristics of local bird populations (Lauha et al., 2022). However, this task is quite challenging as it requires Machine Learning expertise similar to having to train models from scratch. While recent studies have demonstrated progress in applying Machine Learning to accelerate the annotation process in ecoacoustics (Martin et al., 2022; Sethi et al., 2024), these approaches are still under development and not yet widely implemented.
@@
To address these limitations, we have been inspired by methodologies used in camera trap projects (Beery et al., 2019; Rigoudy et al., 2023), in which a detector is applied prior to species classification on captured images. Similarly, we propose a two-stage pipeline approach that uses first a generalizable bird vocalization detector combined later with a classifier, improving accuracy and reducing false positives. In particular, we have developed a Bird Song Detector based on the YOLOv8 model (You Only Look Once v8; Jocher et al. (2023)) trained with data from our case study in Doñana National Park. Once the audio segments containing bird songs have been detected, we applied several classifier models, and in all of them, we have proven an improvement of the classification when the detector was present (Figure 1).
@@ 2.1.1 Field study site and PAM design
Soundscapes were recorded at Doñana National Park (SW Spain). This area corresponds to the marshes of the Guadalquivir delta and is one of the most important wetlands in Southern Europe, where millions of migrating birds stopover and winter every year (Rendón et al., 2008; Green et al., 2016). The area has serious conservation threats due to water over-exploitation for agriculture and tourism, climate and land-use change, and pollution due to a mine spillover in 1998, among others (Green et al., 2024), which poses serious concerns both at the local and European administration levels. These threats are particularly affecting bird populations, which are in decline over the last years (Camacho et al., 2022; Campo-Celada et al., 2022). Doñana includes four major habitats, which are differentiated by their flooding regime and vegetation: coastal dunes, scrublands, marshlands, and the ecotone or transition among them. The deployment design of the BIRDeep project included nine AudioMoth recorders (Hill et al., 2019) that were distributed among three of these habitats: two in the marshland, three in the ecotone and four in the scrubland, differentiating high and low scrubland (see Figure 2). AudioMoths are low-cost automatic audio recording devices with open-source hardware (Hill et al., 2018). They continuously recorded 1 minute of audio every 10 minutes. Configuration parameters of deployed AudioMoth included a sampling rate of 32 kHz, a medium gain, and a filter band focused on bird frequencies (0.6-16.0 kHz).
@@ 2.1.2 Acoustic data annotation by experts
Although devices have been continuously recording since their deployment, for this study, we selected a manually annotated subset spanning from March to October 2023 for logistic reasons. From the total of audios, we selected 461 minutes. We tried to balance representation across sites and habitats while keeping the annotation effort manageable. We also prioritized periods of high bird activity, primarily the morning chorus, to maximize the number of detectable vocalizations in the annotations (Robbins, 1981).
Annotation was carried out manually by two co-authors with ornithological expertise (ESG and GB). The process is labor-intensive and time-consuming, requiring simultaneous listening to the audio and visualization of the spectrogram, which displays how the sound’s frequency content evolves over time. Annotating these 461 minutes required approximately 13,445 minutes (about 224 hours) of labor effort, with a median annotation time of 18 minutes and an average of 29.4 minutes per 1-minute file. The experts used Audacity software (Audacity, 2017), which facilitates species identification using auditory signals and visual patterns based on spectrograms. These annotations were then exported from the Audacity format and converted to CSV files. Each annotation consisted of bounding boxes with the minimum and maximum frequency, as well as the start and end time of each bird vocalization in the spectrogram.
@@
The number of annotations was highly unbalanced across recorders, whereas the number of annotated files was more similar, except for Juncabalejo site (Figure 3). This was the most isolated site in the marshland and we found logistic problems to access the recorders to change batteries and SIM cards during the flooding time. Some recorders had a high number of annotations because their recordings were rich in bird vocalizations, likely reflecting variations in bird occurrence and activity, environmental conditions, and recording quality across sampling sites.
@@
Figure 3: Number of annotations and audio files across sampling sites (see Figure 2 for reference of site names), illustrating the varying levels of annotation effort across recorders.
@@
Figure 4: Number of annotations of the 34 annotated classes in the dataset. All annotated classes, including genus- and family-level labels (e.g., Sturnus sp., Fringillidae) and the general Bird label, were used to train the Bird Song Detector by relabeling them under a unified Bird class. For species classification, only the 29 classes shown in blue were retained. The orange bars represent general or ambiguous classes (e.g., family-level or uncertain identifications) that were excluded from classifier training to ensure specificity.
@@ 2.1.4 Data Preparation
Spectrograms were generated using the librosa Python library (McFee et al., 2015), a popular toolkit for audio analysis. The Short-Time Fourier Transform (STFT) was applied to the audio signal to compute the spectrogram, and the amplitude of the resulting frequency bins was converted to a decibel (dB) scale. The frequency axis was displayed on a logarithmic scale. The frequency range of the spectrograms was set to a minimum of 1 Hz and a maximum of 16,000 Hz, which encompasses the typical range of bird vocalizations while excluding inaudible frequencies or irrelevant noise. The dimensions of the output spectrogram images were $930\times 462$ pixels. Although our annotations originally included frequency windows for each vocalization, to simplify detection given the limited data, bounding boxes spanned the entire frequency y-axis. This choice prioritizes the temporal localization of bird vocalizations, which are the primary signal for subsequent classification (see Figure 6 as an example).
@@
To enhance the dataset and improve the robustness of downstream models, additional preprocessing steps were applied. Specifically, synthetic audio samples were generated by adding random noise (using NumPy Python library (Harris et al., 2020), to simulate background interference and slightly adjust the intensity of the signals. These modifications aimed to increase the number of training examples that contained only background noise, ensuring a more diverse dataset for both detection and classification tasks.
The annotations were further processed to adopt the format required by the YOLOv8-based object detection model (Jocher et al., 2023). A detailed explanation of the mathematical conversion from YOLOv8’s normalized coordinates to temporal annotations in seconds is presented in A.
@@ 2.2.1.2   BirdNET as a bird vocalization detector
    Confidence Score = 0.1 (Default Value): This is the default BirdNET threshold. Prior studies have shown that using this threshold in real-world PAM scenarios leads to an increase in FPs, as BirdNET assigns low-confidence scores to background noise or ambiguous detections (Pérez-Granados, 2023). However, lowering the confidence threshold improves recall by reducing missed detections.
@@
    Confidence Score = 0.6 (Adjusted Threshold): Higher confidence thresholds have been recommended for improving the reliability of BirdNET predictions in complex acoustic environments. Funosas et al. (2024) found that BirdNET predictions become significantly more reliable when using thresholds above 0.7. We adopted a slightly lower threshold of 0.6 to balance precision and recall, reducing FPs while still capturing a sufficient number of detections.
@@ 2.3.2 Other classifiers
During evaluation, the models were tested using both full 1-minute audio files (segmented into 3-second windows) and shorter audio segments identified by the Bird Song Detector. When these detected segments were shorter than the required 3-second input length, they were padded with zeros—representing silence—to match the model’s input size.
@@ 3.2 Selection of confidence score threshold for the Bird Song Detector
On the other hand, higher thresholds, like 80% or 95%, lead to excessive losses of TPs (74.35% and 99.03%, respectively). These thresholds result in an impractical trade-off, where the reduction in FPs comes at the cost of almost complete loss of the TPs (which need to have a high probability threshold to be retained), significantly degrading the detector’s performance.
@@ 3.3 Bird Song Detector Selection
In the case of the Full Frequencies model, although higher precision is achieved, the limited number of predictions raises concerns about its robustness and generalizability. In contrast, the Best Model, with a greater number of predictions and balanced performance metrics, stands out as the more reliable choice for our ecological monitoring objectives, given its consistent performance across metrics.
@@ 4.2 Bird Song Detector vs BirdNET as a Bird Vocalization Detector
Table 2: Comparison of confusion matrices for different models used as bird vocalization detectors, including various configurations of BirdNET (fine-tuned and non fine-tuned, with different confidence thresholds and species lists) and Bird Song Detector model. All models are evaluated solely in terms of detection performance, that is, their ability to distinguish between segments with bird vocalizations and background noise, independently of species classification. The table shows the number of TPs, FPs and FNs for each configuration. The last two rows highlight relative changes in performance when comparing the Bird Song Detector to fine-tuned BirdNET at confidence thresholds of 0.6 (blue) and 0.1 (orange). Bold values indicate improvements made by Bird Song Detector.
@@ 4.3 Bird Species Classification
Table 3: Comparison of classifier performance with and without the Bird Song Detector. Shown metrics are Accuracy (Acc.), macro and weighted average (Avg) precision (Prec.), recall (Rec.), F1-score (F1), and an index calculated based on the number of predictions relative to the total number of annotations (Idx Pred/Annn). Shaded rows indicate improvement when using Bird Song Detector. All metrics are better at higher values, except for Idx Pred/Ann, which is optimal when closer to 1.